Q3 (2 points): Suppose that X takes values between 0 and 1 and has probability density function f(x) = 2x. Compute Var(x) andPlease help with Q3,4,5 and provide steps. Will rate.
Answer #1

Solukon tra- 2x ornal Em = 1! uten = Jo! 2x² = 2/4) EAN) = F142) = lot? 46:11 04 = 1! 2 vºdu = (24). E(35= V() = E (012) - E124 XN N(3,4) M Plx>4) - | - PIX<4) = 1 – Plz24-3) = 1-P(Z <0.5) {z-standard Hommal table } 1 - 0.69146 PIX> +) = 0.3005 PLO <1 +6 - 4 - 경 + = 1thank you